OpenStudy (anonymous):

Can anyone help me solve the following trig question? Show that sin(40)+sin(20)=cos(10)

OpenStudy (anonymous):

use this identity \[\sin p+\sin q=2 \sin \frac{p+q}{2} \ \cos \frac{p-q}{2}\]

OpenStudy (anonymous):

Hmm I got 2sin(30)cos(10) But I still got the 2 sin(30)

OpenStudy (anonymous):

well \(\sin 30=\frac{1}{2}\)

OpenStudy (anonymous):

Oh I see I overlooked that thanks.

OpenStudy (anonymous):

yw :)